# Pressure and strain effects on the ab initio electronic structure of La3Ni2O7

### v1

@@ IV Conclusions
In conclusion, we have studied pressure and biaxial strain effects on the electronic structure of La3Ni2O7, taking into account correlation effects within the ab initio approximation to the self-energy. We confirm that the electronic structure of La3Ni2O7 is extremely sensitive to the internal atomic positions of its crystal structure. correlation effects shift downward the lower Ni-3 band, so to remove the hole pocket from the Fermi level without introducing any adjustable parameter. Therefore, the Fermi surface is composed only of the cuprate-shape sheet plus the nickelate-specific cylinder , both showing an effective 1D character, and is in good agreement with ARPES experiments [56] which did not find any pocket. Additionally, we observe a non-negligible drop in the quasiparticle spectral weight on the Ni-3 states, both and , so to configure a little departure from a Fermi liquid behavior. Finally, we have shown that not only correlations, but also pressure or biaxial strain play a crucial role in pulling down the La-5 band at and M towards the Fermi level, which may induce self-doping. In any case, at pressures and strains of interest of superconductivity, this band might play an important role at low energy, a fact which is often overlooked in effective-model studies.
@@ V Acknowledgments
We thank Markus Holzmann for useful discussions. We acknowledge financing from the LABEX LANEF project NICOS. Computing time has been provided by French GENCI and Grenoble CIGRID.
Appendix A Crystal structures used in the calculations
TAB.Β 1 reports the crystal structures in the conventional cell used for the calculation at 29.5Β GPa and for the in-plane strained system.
TAB.Β 2 reports the crystal structures used for the extra pressure calculations, all obtained via structural optimizations.
| 29.5 GPa [1, 24] | a=3.7148 | c=19.7340 | ||
| Ni | 0 | 0 | 0.096 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.095 | |
| O2 | 0 | 0 | 0.204 | |
| O3 | 0 | 0 | 0 | |
| epitaxial strained (thin film at ambient pressure) [5] | a=3.77 | c=20.6927 | ||
| Ni | 0 | 0 | 0.095 | |
| La1 | 0 | 0 | 0.321 | |
| La2 | 0 | 0 | 0.5 | |
| O1 | 0 | 0.5 | 0.096 | |
| O2 | 0 | 0 | 0.205 | |
| O3 | 0 | 0 | 0 | |
@@ Appendix B Sensitiveness of the electronic structure to internal atomic positions
In our first attempt to calculate the DFT and the electronic structures of the bilayer at 29.5 GPa, we used the experimental atomic positions determined by X-ray diffraction (XRD) given in the extended data table I in Sun et al.Β [1] for the Fmmm phase and adapted to the as also done in Sakakibara et al.Β [24]
The resulting DFT and Fermi surfaces (Fig.Β 6) show the appearance of a cylindrical Fermi sheet centered on which is not observed in the ARPES experiment of Yang et al.Β [56] This fake Fermi sheet originates from the crossing at the Fermi level of the La-5 character band, as one can see in Fig.Β 7a for PBE and in Fig.Β 7b for . It has to be considered as an artifact due to a non correct configuration of the internal atomic positions. In XRD the accuracy on the internal atomic positions is in general lower than on the lattice parameters, especially concerning the oxygen atoms. Due to this drawback we decided for this work to rely on theoretical DFT-PBE relaxed internal atomic positions (see Tab. 1), as explained in Sec. II.2 of the main text.
The DFT bandplot of this relaxed structure is shown on Fig.Β 7c and has to be compared with Fig.Β 7a corresponding to the experimental structure. In particular, we remark that the self-doping induced by the La-5 band is absent for the relaxed structure. The Ni- bands are less affected by these internal atomic positions differences, but quantitative discrepancies are visible, for example the size of the Fermi sheet. Some of these effects were already previously reported [45].
Since any real or fake modification of the electronic structure at the Fermi level may have important consequences for superconductivity, we tried to identify what precisely triggers the La-5 band lowering. To this purpose we computed (at the DFT level only) the bandplot of a structure similar to the relaxed one, but with the La atoms in the spacer (La1 in Tab. 1) artificially shifted away from the NiO2 plane by 0.14Β Γ , which corresponds to 0.7% of (see Fig. 8). The associated bandplot is shown on Fig. 7d and reproduces well the self-doping behavior of the rare-earth band that is observed on the DFT bandplot of Fig.Β 7a, and to a lesser extent on the bandplot of Fig.Β 1d. We could have rather moved the apical oxygens of the spacer toward the NiO2 planes, but this would have changed the shape of the octahedron and potentially the crystal field splitting. In order to keep the ordering of the Ni bands, we have preferred to move the lanthanum of the spacer towards the apical oxygens, which should be equivalent.
@@
We anticipate that and we are working on a more complete study [74] about this effect and its correct description, both the internal atomic positions and the numerical methods employed being important. This can further help shading light on the aspects which could be relevant to superconductivity in RP nickelates.
@@
Appendix C Fermi energy calculation: standard bisection vs DOS integration methods
@@
The indirect gap of 30 meV between the two incipient bands found in the calculation done at 29.5 GPa is beyond the accuracy by which the Fermi level can be calculated. This is already a consequence of the finite smearing temperature used to calculated metals, of the order of Β Ha Β meV and up to Β Ha Β meV (the Abinit default is Β Ha, while it is half that in Quantum Espresso). We must also keep in mind that any change in the parameters used to generate the pseudo-potentials can affect DFT Kohn-Sham energies by even 100200Β meV. This is also the reason why the quasiparticle energy convergence is never pushed beyond 100Β meV. A further indetermination of the Fermi energy arise in particular from the choice of the method used to calculate it. In this Section we present the comparison between two of them: the classical method using a bisection algorithm, as implemented in Abinit or Quantum Espresso; and the Wannier interpolation as implemented in Wannier90, followed by the integration of the total DOS done with our own Python code using the Simpson method.
@@
In this Section we show how critical the calculation of the Fermi energy can be on the electronic structure and the Fermi surface of this system. For this reason, we take the most critical case of 29.5Β GPa and we check the two methods here above described (for short labeled the bisection and DOS methods), keeping as close as possible to their defaults when there is matter of choice. In Fig.Β 9 we show a zoom around the Fermi level of the bandplot, as in Fig.Β 1d but reporting the different Fermi energies calculated using a smearing temperature of Ha, with the corresponding Fermi surfaces below. First of all, one can remark the robustness of the and sheets with respect to a large variation of (120 meV), which is due to the more dispersive character of these bands compared with to two incipient ones.
@@
With the bisection method ( traced by a black line and set as the origin), a pocket of electrons () appears around the M point in the plane (see BZ scheme on Fig.Β 1c), due to the La-5 band crossing. On the other hand, with the DOS method and using the least arbitrary Wannier functions set among all possible sets, the PWF one, the Fermi energy is 87 meV (purple line) lower. In this case, there is no pocket of electrons at M and, in contrast, a pocket of holes is opened in the XP direction in the calculation, and the Fermi surface becomes again very similar to the DFT-PBE one, i.e.Β with the reappearance of the pocket in the BZ corners. As it can be seen, different methods provide qualitatively different Fermi surfaces leading to different low energy physics.
@@
In order to improve our estimate on the Fermi energy, we have performed a full DFT+ calculation reducing the smearing temperature to Ha Β meV. The corresponding quasiparticle energies are indicated by pink stars and the associated by the green dashed line in Fig.Β 9. As we can see, now the Fermi level situated more or less in the middle of the indirect band gap, at meV compare to the calculated with Ha. As a consequence both the electron pocket and the hole pocket are suppressed. And this is the Fermi level we have used to plot the Fermi surface in Fig.Β 1b.
@@
In practice, in a system with such a delicate balance, any defect, vacancy (La3Ni2O7 is usually oxygen deficient) or impurity, introduces a tiny perturbation in the Fermi energy which will open the electron or the hole pocket. In any case, the Fermi surface showed on Fig.Β 1b corresponds to a theoretical ideal situation. Therefore, we expect that a real experiment will probably not measure the ideal situation. Therefore we leave a margin of error of 50Β meV on the Fermi level to keep into account both experimental and theoretical uncertainties.
@@
Notice that 29.5Β GPa is a really critical pressure for the La3Ni2O7 electronic structure. Indeed, at ambient pressure the La-5 band is almost 0.5Β eV above its position at 29.5Β GPa (see Fig.Β 4b), and the bisection method set the Fermi energy well below the bottom of this band, with no possibility of an electron pocket opened at M. Consequently, the Fermi surface would be exactly that in Fig.Β 1b, with the stable and sheets and no or pockets at all, from ambient pressure on, and without any relevant modification up to 29.5Β GPa.
@@
Appendix D Orbital character projection: pure atomic orbitals vs Wannier functions
@@
Fig. 10 shows three DFT PBE band character plots of La3Ni2O7 at 29.5 GPa: the first one using projections on (pseudo) atomic orbitals calculated with Quantum Espresso, which we will consider as the reference; and the two others with Wannier90 but using different WF sets, namely projected Wannier functions (PWF) and maximally-localized Wannier functions (MLWF). PWF are the Wannier90 first guess and correspond to WF with imposed well defined orbital character (in the input file), though not purely hydrogenoid atomic orbitals [75, 76] (see below). Therefore, PWF is the set of WF one obtains from Wannier90 by setting to zero the number of iterations in the Wannierisation procedure (not in the disentanglement one). PWF might not be the set best suited to a smooth interpolation of bands. However, in this case we used 67 WF, which is a set large enough to allows an already smooth interpolation. On the other hand, MLWF are obtained after a number of Wannierisation iterations sufficient to minimize the Marzari-Vanderbilt functional towards the global or a local minimum, which correspond to a maximum in the localization of the WF. MLWF can loose their original orbital character, and they can even localize outside atoms and in the interstitial, if the target is to have the most localized WF.
@@
The Quantum Espresso and the PWF band plot characters are very consistent between themselves, whereas the MLWF shows important discrepancies for the La- orbital character which lose weight on the band of interest and is more hybridized with the 4 manifold. It must be kept into account that the projection onto MLWF, beyond to strongly depend on the chosen set of MLWF (e.g.Β localized on atoms, on bonds, or on anti-bonding positions), is different from the projection onto atomic wave functions. But there is large arbitrariness also for atomic wave functions. For example whether projecting onto all-electron wave functions (as in all-electron codes), or pseudo-atomic wave functions (e.g.Β Quantum Espresso and Abinit, with some minor differences between them), or spherical harmonics (e.g.Β VASP). In conclusion, projections can be a convenient label to identify bands, establishing a convention. However, any physics argument built on them is always risky.
@@
Appendix E Additional orbital character projections
On Fig.Β 12 we provide more precise band character plots than on Fig.Β 2b, giving details on the specific orbitals involved in the Bloch states. We distinguish the two spacer La atoms from the inner one lying in between the two NiO2 planes; as well as the planar O atoms lying in the NiO2 planes, and the apical O atoms of the spacer (ap.Β sp.) or of the inner layer (ap.Β inner). See Fig.Β 13 which shows the atoms labeling on the Vesta structure. For O-2, we also distinguish and orbitals.
We remark that there is a non-negligible La-4 orbital character on the lowered La-5 band which almost reaches the Fermi level. We also remark that in-plane O-2 orbitals have a strong contribution on what we label Ni-3 bands, especially above the Fermi level.
Appendix F Convergence in the number of bands
In a calculation, the number of bands taken into account in the calculation of the self-energy is critical. A convergence study is therefore needed. Moreover, this study is motivated by the discrepancy on the Ni-3 upper and lower bands which we have found between our results and the results of Christiansson et al.Β [45] from one side, and of You et al.Β [15] from another, especially regarding the shift the latter obtained on the upper Ni-3 band from DFT to . In our original calculation, we used 225 bands. For this test we rise up to 300 for the calculation of energies of three bands around the Fermi level. The energies obtained are plotted on top of the original calculation in Fig.Β 14. As we can see, the original calculation is already enough converged, the maximal difference for the set of checked bands being of Β meV, validating the electronic structure presented in the main text.
Appendix G Fermi surfaces of the substrate constrained and ambient pressure cases
Appendix H Bloch state plots

### v2

@@ IV Conclusions
In conclusion, we have studied pressure and biaxial strain effects on the electronic structure of La3Ni2O7, taking into account correlation effects within the ab initio approximation to the self-energy. We confirm that the electronic structure of La3Ni2O7 is extremely sensitive to the internal atomic positions of its crystal structure. correlation effects shift downward the lower Ni-3 band, so to remove the hole pocket from the Fermi level without introducing any adjustable parameter. Therefore, the Fermi surface is composed only of the cuprate-shape sheet plus the nickelate-specific cylinder , both showing an effective 1D character, and is in good agreement with ARPES experiments [56] which did not find any pocket. Additionally, we observe a nonnegligible drop in the quasiparticle spectral weight on the Ni-3 states, both and , so to configure a little departure from a Fermi liquid behavior. Finally, we have shown that not only correlations, but also pressure or biaxial strain play a crucial role in pulling down the La-5 band at and M towards the Fermi level, which may induce self-doping. In any case, at pressures and strains of interest of superconductivity, this band might play an important role at low energy, a fact which is often overlooked in effective-model studies.
@@ V Acknowledgments
We thank Markus Holzmann for useful discussions. We acknowledge financing from the LABEX LANEF project NICOS. Computing time has been provided by French GENCI Grant No. 2022-AD01091394 and Grenoble CIGRID.
| 29.5 GPa [1, 24] | a=3.7148 | c=19.7340 | ||||
| Ni | 0 | 0 | 0.096 | |||
| La1 | 0 | 0 | 0.321 | |||
| La2 | 0 | 0 | 0.5 | |||
| O1 | 0 | 0.5 | 0.095 | |||
| O2 | 0 | 0 | 0.204 | |||
| O3 | 0 | 0 | 0 | |||
a=3.77 | c=20.6927 | ||||
| Ni | 0 | 0 | 0.095 | |||
| La1 | 0 | 0 | 0.321 | |||
| La2 | 0 | 0 | 0.5 | |||
| O1 | 0 | 0.5 | 0.096 | |||
| O2 | 0 | 0 | 0.205 | |||
| O3 | 0 | 0 | 0 | |||
@@
Appendix A Crystal structures used in the calculations
TAB. 1 reports the crystal structures in the conventional cell used for the calculation at 29.5 GPa and for the in-plane strained system.
TAB. 2 reports the crystal structures used for the extra pressure calculations, all obtained via structural optimizations.
@@ Appendix B Sensitiveness of the electronic structure to internal atomic positions
In our first attempt to calculate the DFT and the electronic structures of the bilayer at 29.5 GPa, we used the experimental atomic positions determined by X-ray diffraction (XRD) given in the extended data table I in Sun et al. [1] for the Fmmm phase and adapted to the as also done in Sakakibara et al. [24]
The resulting DFT and Fermi surfaces (Fig. 6) show the appearance of a cylindrical Fermi sheet centered on which is not observed in the ARPES experiment of Yang et al. [56] This fake Fermi sheet originates from the crossing at the Fermi level of the La-5 character band, as one can see in Fig. 7a for PBE and in Fig. 7b for . It has to be considered as an artifact due to a non correct configuration of the internal atomic positions. In XRD the accuracy on the internal atomic positions is in general lower than on the lattice parameters, especially concerning the oxygen atoms. Due to this drawback we decided for this work to rely on theoretical DFT-PBE relaxed internal atomic positions (see Tab. 1), as explained in Sec. II.2 of the main text.
The DFT bandplot of this relaxed structure is shown on Fig. 7c and has to be compared with Fig. 7a corresponding to the experimental structure. In particular, we remark that the self-doping induced by the La-5 band is absent for the relaxed structure. The Ni- bands are less affected by these internal atomic positions differences, but quantitative discrepancies are visible, for example the size of the Fermi sheet. Some of these effects were already previously reported [45].
Since any real or fake modification of the electronic structure at the Fermi level may have important consequences for superconductivity, we tried to identify what precisely triggers the La-5 band lowering. To this purpose we computed (at the DFT level only) the bandplot of a structure similar to the relaxed one, but with the La atoms in the spacer (La1 in Tab. 1) artificially shifted away from the NiO2 plane by 0.14 Å, which corresponds to 0.7% of (see Fig. 8). The associated bandplot is shown on Fig. 7d and reproduces well the self-doping behavior of the rare-earth band that is observed on the DFT bandplot of Fig. 7a, and to a lesser extent on the bandplot of Fig. 1d. We could have rather moved the apical oxygens of the spacer toward the NiO2 planes, but this would have changed the shape of the octahedron and potentially the crystal field splitting. In order to keep the ordering of the Ni bands, we have preferred to move the lanthanum of the spacer towards the apical oxygens, which should be equivalent.
@@
Appendix C vs band plots
In this Section we report the band plot for the Amam crystal structure relaxed at GPa, shown in its oS base-centered orthorhombic BZ (Fig. 9 red lines). For comparison we also show the band plot for the 0 GPa relaxed system but constrained within the I4/mmm crystal structure and folded within the same oS BZ (blue lines) from its original tI body-centered tetragonal (BCT) BZ. Of course, the two BZ have different topology and even after the folding they do not overlap. High symmetry points of the tI may not correspond to any high symmetry point of the oS. And even in case of one-to-one correspondence between high symmetry points, they might be located at slightly different positions due to the different BZ shape determined by the to (and to ) ratios which of course differ.
We remark some differences, especially in the conduction states which however are not observed by direct ARPES photoemission, due to the reduction of symmetry induced by the tilting of NiO6 octahedra and by the introduced orthorhombicity which are both at their maximum at 0 GPa. Nevertheless, the two plots are very similar at the Fermi level. In particular the presence of the hole pocket is confirmed also in the DFT PBE Amam structure: only folded from the X-P segment of the BCT Brillouin zone, into the (initial part of the) -Y segment of the oS BZ. The only noticeable difference is, in the Amam structure, a slightly larger dispersion of the band along the axis than in the I4/mmm where it is almost undispersive. The Amam band is either at the same level (in ) or shifted meV downward (in Y) compared to the I4/mmm one. We know that the general tendency of the corrections is to shift down this band by meV with respect to DFT PBE. Therefore the pocket should disappear also in a calculation of the Amam structure.
Finally, outside the Fermi region we remark that the La-5 band is further shifted up in the Amam crystal structure with respect to the I4/mmm where, as discussed in the main text, it is found a little below 1 eV at the M point (corresponding to the Y point of the oS). Therefore this band is further outside the low-energy region for the combined effect of pressure and symmetry reduction, and enter into play at the pressure of 14 GPa, as illustrated in Fig. 4b.
This comparison indicates that the main conclusions of this work, and in particular about the topology of the Fermi surface and the disappearance of the hole pocket, should also hold for the case of the Amam crystal structure at ambient pressure as measured by ARPES.
@@
Appendix D Fermi energy calculation: standard bisection vs DOS integration methods
@@
The indirect gap of 30 meV between the two incipient bands found in the calculation done at 29.5 GPa is beyond the accuracy by which the Fermi level can be calculated. This is already a consequence of the finite smearing temperature used to calculated metals, of the order of Ha meV and up to Ha meV (the Abinit default is Ha, while it is half that in Quantum Espresso). We must also keep in mind that any change in the parameters used to generate the pseudopotentials can affect DFT Kohn-Sham energies by even 100200 meV. This is also the reason why the quasiparticle energy convergence is never pushed beyond 100 meV. A further indetermination of the Fermi energy arise in particular from the choice of the method used to calculate it. In this Section we present the comparison between two of them: the classical method using a bisection algorithm, as implemented in Abinit or Quantum Espresso; and the integration of the Wannier90 total DOS done with our own Python code using the Simpson method. The Wannier90 DOS is calculated from the PWF-interpolation bands, on a grid.
@@
In this Section we show how critical the calculation of the Fermi energy can be on the electronic structure and the Fermi surface of this system. For this reason, we take the most critical case of 29.5 GPa and we check the two methods here above described (for short labeled the bisection and DOS methods), keeping as close as possible to their defaults when there is matter of choice. In Fig. 10 we show a zoom around the Fermi level of the bandplot calculated using a smearing temperature of Ha, as in Fig. 1d, but reporting the different Fermi energies calculated and the associated Fermi surfaces below. First of all, one can remark the robustness of the and sheets with respect to a large variation of (100 meV), which is due to the more dispersive character of the corresponding bands compared with to two incipient ones.
@@
With the bisection method ( traced by a black line and set as the origin), a pocket of electrons () appears around the M point in the plane (see BZ scheme on Fig. 1c), due to the La-5 band crossing. On the other hand, with the DOS method and using the least arbitrary Wannier functions set among all possible sets, the PWF one, the Fermi energy is 87 meV (purple line) lower. In this case, there is no pocket of electrons at M and, in contrast, a pocket of holes is opened in the XP direction in the calculation, and the Fermi surface becomes again very similar to the DFT-PBE one, i.e. with the reappearance of the pocket in the BZ corners. As it can be seen, different methods provide qualitatively different Fermi surfaces leading to different low energy physics.
@@
In order to improve our estimate on the Fermi energy, we have performed a full DFT+ calculation reducing the smearing temperature to Ha meV. The corresponding quasiparticle energies are indicated by green stars and the associated by the green dashed line in Fig. 10. As we can see, now the Fermi level situated more or less in the middle of the indirect band gap, at meV compare to the calculated with Ha. As a consequence both the electron pocket and the hole pocket are suppressed. And this is the Fermi level we have used to plot the Fermi surface in Fig. 1b.
@@
In practice, in a system with such a delicate balance, any defect, vacancy (La3Ni2O7 is usually oxygen deficient) or impurity, introduces a tiny perturbation in the Fermi energy which will open the electron or the hole pocket. In any case, the Fermi surface showed on Fig. 1b corresponds to a theoretical ideal situation. Therefore, we expect that a real experiment will probably not measure the ideal situation, and we leave a margin of error of 100 meV on the Fermi level to keep into account both experimental and theoretical uncertainties.
@@
Notice that 29.5 GPa is a really critical pressure for the La3Ni2O7 electronic structure. Indeed, at ambient pressure the La-5 band is almost 0.5 eV above its position at 29.5 GPa (see Fig. 4b), and the bisection method set the Fermi energy well below the bottom of this band, with no possibility of an electron pocket opened at M. Consequently, the Fermi surface would be exactly that in Fig. 1b, with the stable and sheets and no or pockets at all, from ambient pressure on, and without any relevant modification up to 29.5 GPa.
@@
Appendix E Orbital character projection: pure atomic orbitals vs Wannier functions
@@
Fig. 11 shows three DFT PBE band character plots of La3Ni2O7 at 29.5 GPa: the first one using projections on (pseudo) atomic orbitals calculated with Quantum Espresso, which we will consider as the reference; and the two others with Wannier90 but using different WF sets, namely projected Wannier functions (PWF) and maximally localized Wannier functions (MLWF). PWF are the Wannier90 first guess and correspond to WF with imposed well defined orbital character (in the input file), though not purely hydrogenoid atomic orbitals [74, 75] (see below). Therefore, PWF is the set of WF one obtains from Wannier90 by setting to zero the number of iterations in the Wannierisation procedure (not in the disentanglement one). PWF might not be the set best suited to a smooth interpolation of bands. However, in this case we used 67 WF, which is a set large enough to allows an already smooth interpolation. On the other hand, MLWF are obtained after a number of Wannierisation iterations sufficient to minimize the Marzari-Vanderbilt functional towards the global or a local minimum, which correspond to a maximum in the localization of the WF. MLWF can loose their original orbital character, and they can even localize outside atoms and in the interstitial, if the target is to have the most localized WF.
@@
The Quantum Espresso and the PWF band plot characters are very consistent between themselves, whereas the MLWF shows important discrepancies for the La- orbital character which lose weight on the band of interest and is more hybridized with the 4 manifold. It must be kept into account that the projection onto MLWF, beyond to strongly depend on the chosen set of MLWF (e.g. localized on atoms, on bonds, or on antibonding positions), is different from the projection onto atomic wave functions. But there is large arbitrariness also for atomic wave functions. For example whether projecting onto all-electron wave functions (as in all-electron codes), or pseudoatomic wave functions (e.g. Quantum Espresso and Abinit, with some minor differences between them), or spherical harmonics (e.g. VASP). In conclusion, projections can be a convenient label to identify bands, establishing a convention. However, any physics argument built on them is always risky.
@@
One can remark on Fig. 12 that PWF do not correspond to real atomic orbitals, showing extra lobes and adaptation to neighboring atoms. This is due to the projection onto Bloch states and to the Löwdin orthogonalization procedure [75] as described in Souza et al. [74]
Appendix F Additional orbital character projections
On Fig. 15 we provide more precise band character plots than on Fig. 2b, giving details on the specific orbitals involved in the Bloch states. We distinguish the two spacer La atoms from the inner one lying in between the two NiO2 planes; as well as the planar O atoms lying in the NiO2 planes, and the apical O atoms of the spacer (ap. sp.) or of the inner layer (ap. inner). See Fig. 13 which shows the atoms labeling on the Vesta structure. For O-2, we also distinguish and orbitals.
We remark that there is a nonnegligible La-4 orbital character on the lowered La-5 band which almost reaches the Fermi level. We also remark that in-plane O-2 orbitals have a strong contribution on what we label Ni-3 bands, especially above the Fermi level.
Appendix G Convergence in the number of bands
In a calculation, the number of bands taken into account in the calculation of the self-energy is critical. A convergence study is therefore needed. Moreover, this study is motivated by the discrepancy on the Ni-3 upper and lower bands which we have found between our results and the results of Christiansson et al. [45] from one side, and of You et al. [15] from another, especially regarding the shift the latter obtained on the upper Ni-3 band from DFT to . In our original calculation, we used 225 bands. For this test we rise up to 300 for the calculation of energies of three bands around the Fermi level. The energies obtained are plotted on top of the original calculation in Fig. 14. As we can see, the original calculation is already enough converged, the maximal difference for the set of checked bands being of meV, validating the electronic structure presented in the main text.
Appendix H Fermi surfaces of the substrate constrained and ambient pressure cases
Appendix I Bloch state plots